If a variable straight line xcos¤ +ysin¤ = p (p is a constant) is a chord of a hyperbola x^2/a^2 + y^2/b^2 = 1 (b>a), substends a right angle at the centre of the hyperbola, then it always touches a fixed circle whose radius is:a) ab/(b-2a)^(1/2)b) a/ (a-b)^(1/2)c) ab/ (b^2 - a^2)^(1/2)d) ab/b (b+a)^(1/2)

Grade:11

1 Answers

Shreyansh Shukla
4 years ago
Equation of pair of straight lines passing through the origin(centre of hyperbola) and points of inresection of the variable chord & hyperbola is:
\frac{x^2}{a^2}-\frac{y^2}{b^2}-(\frac{xcos\alpha+ysin\alpha }{p})^2=0
They are at right angles if coefficient of x^2coefficient of y^2=0 i.e.
 
(\frac{1}{a^2}-\frac{(cos\alpha)^2 }{p^2})-(\frac{1}{b^2}+\frac{(sin\alpha )^2}{p^2})=0
 
\frac{1}{a^2}-\frac{1}{b^2}=\frac{1}{p^2}
 
p=\frac{ab}{\sqrt{b^2-a^2}}
 
As p is length of perpendicular form origin on the line xcos\alpha +ysin\alpha =p, the line touches the circle with centre at origin and radius = \frac{ab}{\sqrt{b^2-a^2}}